OpenStudy (anonymous):

if (3x^2)+7=0 then (x-1/3)^2 how do you suppose to solve this problem?

OpenStudy (anonymous):

is the 3 under x-1? or under 1 only?

OpenStudy (anonymous):

x^2=-7/3 ..... (1) (x-1/3)^2=x^2-2/3x+1/9= from (1) we get \[-7/3-2/3\sqrt{7/3}i +1/9\] \[=-20/9-2/3\sqrt{7/3}i\] whic i=imajiner